# SI_appendix_phys

### v1

@@ 3 Traveling solutions in large systems
The nonlinear systems outer0,outer1,outer2 are again solved numerically using NonlinearSolve() with paramters reltol=1e-8, abstol=1e-8 and maxiters=20. We first initialise outer0 using the finite domain solution (Subsection 2) with $\phi=0.80,\phi_{a}=0.45$, $\text{Pe}=7.5$, $L=100$, $N=3200$. Once an initial solution to outer0 has been found, we initialize the iterative solver using an existing solution with similar parameters: $\phi_{a}^{\prime}=\phi_{a}\pm 0.01$, $\phi_{p}^{\prime}=\phi_{p}\pm 0.01$. We first initialise outer1 using the finite domain solution (Subsection 2) with $\phi=0.8,\phi_{a}=0.3$,$\text{Pe}=7.5$,$N=3200$. Once an initial solution to outer0 has been found, we initialize the iterative solver using an existing solution with similar parameters: $\phi^{\prime}=\phi\pm 0.01$, $\nu^{\prime}=\nu\pm 0.01$. We first initialise outer2 using the outer1 solution with $\phi=0.65,\nu=0.36$,$\text{Pe}=7.5$,$N=1024$, and place the new interface at $j=299$. Once an initial solution to outer2 has been found, we initialize the iterative solver using an existing solution with similar parameters: $\nu^{\prime}=\nu\pm 0.01$, $j^{\prime}=j\pm 8$.
Figure S4: Profiles obtained by solving Eq. (8). (a) T solution. (b) PS solution. The top panels display the solution profile as a function of $z/L$. The bottom panels display the profile (solid black) overlayed on phase diagrams spanned by $\phi_{a}$ and $\phi_{p}$. The Spinodal (Blue) encloses the region of linear stability of homogeneous solutions, and the Binodal (Red) encloses the region of Phase separation. Black crosses mark the volume fraction. Parameters: $\text{Pe}=7.5$, $\phi_{a}=0.4608$, $\phi_{p}=0.2201$, $N=1024$

### v2

@@ 3 Traveling solutions in large systems
The nonlinear systems outer0,outer1,outer2 are again solved numerically using NonlinearSolve() with paramters reltol=1e-8, abstol=1e-8 and maxiters=20. We first initialise outer0 using the finite domain solution (Subsec. 2) with $\phi=0.80,\phi_{a}=0.45$, $\text{Pe}=7.5$, $L=100$, $N=3200$. Once an initial solution to outer0 has been found, we initialize the iterative solver using an existing solution with similar parameters: $\phi_{a}^{\prime}=\phi_{a}\pm 0.01$, $\phi_{p}^{\prime}=\phi_{p}\pm 0.01$. We first initialise outer1 using the finite domain solution (Subsec. 2) with $\phi=0.8,\phi_{a}=0.3$,$\text{Pe}=7.5$,$N=3200$. Once an initial solution to outer0 has been found, we initialize the iterative solver using an existing solution with similar parameters: $\phi^{\prime}=\phi\pm 0.01$, $\nu^{\prime}=\nu\pm 0.01$. We first initialise outer2 using the outer1 solution with $\phi=0.65,\nu=0.36$,$\text{Pe}=7.5$,$N=1024$, and place the new interface at $j=299$. Once an initial solution to outer2 has been found, we initialize the iterative solver using an existing solution with similar parameters: $\nu^{\prime}=\nu\pm 0.01$, $j^{\prime}=j\pm 8$.
Appendix F CP and T solutions
Fig. S5 illustrates a CP solution in a large domain, showing the emergence of narrow interfaces as in T states. Fig. S6 shows an example state point where both T and PS solutions can exist – this is multistability, as discussed in Sec..5 of main text.
To further clarify the relationship between CP and T solutions, we sketch in Fig. S7 the qualitative behavior that we expect in large systems. The T solution involves a localized set of active particles that “push” a set of passive ones through the system (Fig. S7(a)). Red arrows indicate the positive self-propulsive forces that act in the (inner) region where $\rho_{a}$ rises steeply; $m$ is positive here and the relevant terms in the equation of motion are those proportional to Pe. Blue arrows indicate that the passive particles are pushed by the active ones, primarily due to cross-diffusion terms in the equation of motion.
The CP solution consists of two localized sets of active particles moving in opposite directions, resembling a combination of two T solutions traveling in opposite directions (Fig. S7(b)). There are two associated inner regions where these particles propel themselves (red arrows), and the active particles push the passive ones as before. When the two traveling objects approach each other, they collide. The collision process is complex, with the clusters departing the collision at high speeds. (Note: we depict the traveling objects as localized for illustrative purposes, but the actual densities are functions of $x/L$ in the outer region.)
An essential feature of a stable CP solution is that both traveling objects survive the collision. In contrast, while the early-time behavior shown in Fig. 4 resembles a CP solution, after three collisions, it becomes evident that the two traveling objects have unequal masses, resulting in only one object surviving after the fourth collision. The outcome for large times corresponds to a T solution.